Find the weight density of a block 7 in. x 5.5 in. x 9 in. that weighs 65 lb. Express your
answer to four decimal places.
1
Expert's answer
2020-04-06T08:54:24-0400

The weight density 

"\rho=\frac{m}{V}=\frac{65\:\rm lb}{7\times 5.5\times 9\:\rm in^3}=0.1876\:\rm lb/in^3"